9. Forecast the level of Q when P = 5, 10, 20, and 30 given the following information:

Qx = 100, 88, 76, 65, 54, 47, 35, 28, 21, 18, 10, 8, 6, 3, 1

Px = 1, 4, 8, 12, 15, 18, 22, 26, 29, 31, 35, 40, 44, 48
Answer #1

By using R software,

83% | 2:44 PM o rextester.com/l/r_onlin 21 : I rextester Language: R qx=c( 100, 88, 76, 65, 54, 47, 35, 28, 21, 18, 10, 8, 6,3 83% | 2:44 PM 1 . 045 Call: lm( formula = qx - px) Coefficients: (Intercept) 89.948 px -2.103 Call: lm(formula = qx - px) R3 83% | 2:44 PM 1 . Call: lm( formula = qx - px) Residuals: Min 1Q Median 3Q Max -8.684 -6.654 -2.059 5.566 13.993 Coefficien

Now, predicted value for given values 5,10,20,30 are respectively 79.4335,68.9185,47.8885,26.8585